OpenStudy (anonymous):

Write the standard equation for the circle center (–6, –8), that passes through (0, 0).

OpenStudy (anonymous):

\[( x+6)^ + (y+8)^2= r^2 \] What is r?

OpenStudy (anonymous):

\[( x+6)^2 + (y+8)^2= r^2 \]

OpenStudy (anonymous):

Reverse the center coordinates and square the radius. Plug it into (x - h)^2 + (y - k)^2 = r^2

OpenStudy (anonymous):

\[ r=\sqrt{ 6^2 + 8^2} =10 \]

OpenStudy (anonymous):

\[( x+6)^2 + (y+8)^2=100 \]